# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 6. Proofs
could be uniformly bounded from above on $[-C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $[-C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
@@
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Since $11-\bar{\lambda}/M>$ 1, there exists $k^{\prime}\in\mathbb{N}$ such that $\bar{\lambda}\left(1-\bar{\lambda}/{M}\right)^{2k}<R/\nu(\mathcal{Y}\cap\mathcal{O}_{2})$ holds for all $k\geq k^{\prime}$. Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
@@
$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$
$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$
$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$
which completes the proof. ∎
@@
Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS. Then,
$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle\Sigma_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes\Sigma_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes\Sigma_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
@@
$F({\boldsymbol{x}})=\Sigma^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
[…]
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg(\Sigma^{\prime}_{0}),\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}\Sigma^{\prime}\right\}.$
@@
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{\Sigma}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{\Sigma}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (16)
@@
$\overline{\Sigma}_{0}\coloneqq\Sigma^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{\Sigma}\coloneqq\Sigma^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
Since $\Sigma^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}\Sigma^{\prime}$, we have $\overline{\Sigma}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times m}$ for some $\ell\in\mathbb{N}$, are homogeneous in ${\boldsymbol{x}}$ of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively. Thus,
$\overline{\Sigma}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
@@
$\overline{\Sigma}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
@@
Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $\Sigma_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $\Sigma\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion. ∎
@@
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ ($\deg(S_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in ${\boldsymbol{z}}$ (resp., ${\boldsymbol{z}}$ and ${\boldsymbol{y}}$), such that
@@
Consider the polynomial matrices $F({\boldsymbol{z}}\circ{\boldsymbol{z}})\in{\mathbb{S}}[{\boldsymbol{z}}]^{p}$ and $G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{q}$. Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ which are homogeneous in ${\boldsymbol{z}}$, such that
@@
Moreover, by Theorem 4.1, $\deg(H_{0})=N+\deg F$ and $\deg_{{\boldsymbol{x}}}(H)=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
$\{H_{0,{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}:{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
@@
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$. As +${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical. Letting $S_{0}\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{0,{\boldsymbol{\alpha}}}^{\intercal}H_{0,{\boldsymbol{\alpha}}})$ and $S\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}(H_{{\boldsymbol{\alpha}}}^{\intercal}H_{{\boldsymbol{\alpha}}})$, the conclusion follows. ∎
@@
[45, Theorem 3] Suppose that the polynomial matrix $P({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $P({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$. Write $P({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}P_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $P_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
$L(P)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(P)}\frac{{\boldsymbol{\alpha}}!}{\deg(P)!}\|P_{{\boldsymbol{\alpha}}}\|,$
[…]
$N\geq\frac{\deg(P)(\deg(P)-1)L(P)}{2\lambda}-\deg(P),$
all coefficients of $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}P({\boldsymbol{x}})$ are PD.
We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg(F),\deg_{{\boldsymbol{x}}}(G)\right\}$.
@@
$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg(F)}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
@@
By homogeneity, we only need to prove $F_{\varepsilon}\succ 0$ on $\overline{\mathcal{X}}\coloneqq{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}.$ That is, there exists $\varepsilon>0$ such that $F({\boldsymbol{x}})-\varepsilon\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$ for all ${\boldsymbol{x}}\in\overline{\mathcal{X}}$. Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on +${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
@@
the conclusion follows. ∎
We now prove the following Pólya-type Positivstellensatz for polynomial matrices with UQs.

### v4

@@ 6. Proofs
could be uniformly bounded from above on $C,C]^{n}\times\mathcal{Y}$ for all $k\in\mathbb{N}$. Therefore, as $C,C]^{n}$ and $\mathcal{Y}$ are compact, there exists $R\in{\mathbb{R}}$ such that
$F({\boldsymbol{x}})-\sum_{j\neq j_{0}}\int_{\mathcal{Y}}\left\langle I_{p}\otimes\left(I_{q_{j}}-G_{j}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$ (21)
@@
for all ${\boldsymbol{x}}\in[-C,C]^{n}$. Since $1-\bar{\lambda}/M>1$, there exists $k^{\prime}\in\mathbb{N}$ such that
$\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}<\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}$ (22)
holds for all $k\geq k^{\prime}$. Then, for all ${\boldsymbol{x}}\in\mathcal{O}_{1}\cap[-C,C]^{n}$ and $k\geq k^{\prime}$,
@@
$\succeq$ $-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left\langle I_{p}\otimes\left(I_{q_{j_{0}}}-G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})/M\right)^{2k},G_{j_{0}}({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+(Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R)I_{p}$ (by (21))
$\succeq$ $\left(-\int_{\mathcal{Y}\cap\mathcal{O}_{2}}\left(Mq^{\prime}+\bar{\lambda}\left(1-\frac{\bar{\lambda}}{M}\right)^{2k}\right)\mathrm{d}\nu({\boldsymbol{y}})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}$ (by (20))
$\succeq$ $\left(-\left(Mq^{\prime}+\frac{R}{\nu(\mathcal{Y}\cap\mathcal{O}_{2})}\right)\nu(\mathcal{Y}\cap\mathcal{O}_{2})+Mq^{\prime}\nu(\mathcal{Y}\cap\mathcal{O}_{2})+1+R\right)I_{p}=I_{p},$ (by (22))
which validates the claim and completes the proof. ∎
@@
Write $h=\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})$ where all $\sigma_{0}\in{\mathbb{R}}[{\boldsymbol{x}}]$ and $S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q_{j}}$ are SOS. Then
$Sh=S\sigma_{0}+\sum_{j=1}^{s}S\int_{\mathcal{Y}}\langle S_{j},G_{j}\rangle\mathrm{d}\nu({\boldsymbol{y}})=S\sigma_{0}+\sum_{j=1}^{s}\int_{\mathcal{Y}}\left\langle S\otimes S_{j},G_{j}\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
As $S\sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $S\otimes S_{j}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq_{j}}$ are all SOS matrices, we have $Sh\in\mathcal{Q}^{p}(\mathbf{G},\nu)$. ∎
@@
$F({\boldsymbol{x}})=S^{\prime}_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S^{\prime}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})+H({\boldsymbol{x}})(1-\|{\boldsymbol{x}}\|^{2}),$
where $^{\prime}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$^{\prime}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ are SOS matrices and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$. Replacing $\boldsymbol{x}$} by ${\boldsymbol{x}}/\|{\boldsymbol{x}}\|$ in the above equality yields
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{-\deg{F}}=S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)+\int_{\mathcal{Y}}\left\langle S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right),G\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
$k^{\prime}\coloneqq\max\left\{\deg{F},\ \deg S^{\prime}_{0},\ \deg_{{\boldsymbol{x}}}{G}+\deg_{{\boldsymbol{x}}}S^{\prime}\right\}.$
@@
$F({\boldsymbol{x}})\|{\boldsymbol{x}}\|^{k^{\prime}-\deg{F}}=\overline{S}_{0}\left({\boldsymbol{x}}\right)+\int_{\mathcal{Y}}\left\langle\overline{S}\left({\boldsymbol{x}},{\boldsymbol{y}}\right),G\left({\boldsymbol{x}},{\boldsymbol{y}}\right)\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}),$ (23)
@@
$\overline{S}_{0}\coloneqq S^{\prime}_{0}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|}\right)\|{\boldsymbol{x}}\|^{k^{\prime}}\ \ \text{and}\ \ \overline{S}\coloneqq S^{\prime}\left(\frac{{\boldsymbol{x}}}{\|{\boldsymbol{x}}\|},{\boldsymbol{y}}\right)\|{\boldsymbol{x}}\|^{k^{\prime}-\deg_{{\boldsymbol{x}}}{G}}.$
Since $S^{\prime}$ is an SOS matrix and $k^{\prime}-\deg_{{\boldsymbol{x}}}{G}\geq\deg_{{\boldsymbol{x}}}S^{\prime}$, we have $\overline{S}=H^{\intercal}H$ with $H=H_{1}+H_{2}\|{\boldsymbol{x}}\|$ where $H_{1},H_{2}\in{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell\in\mathbb{N}$, are homogeneous in $\boldsymbol{x}$} of degree $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2$ and $(k^{\prime}-\deg_{{\boldsymbol{x}}}{G})/2-1$, respectively. Thus,
$\overline{S}=H^{\intercal}H=(H_{1}+H_{2}\|{\boldsymbol{x}}\|)^{\intercal}(H_{1}+H_{2}\|{\boldsymbol{x}}\|)=\left(H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{1}^{\intercal}H_{2}+H_{2}^{\intercal}H_{1}\right)\|{\boldsymbol{x}}\|.$
@@
$\overline{S}_{0}=\left(H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}\right)+\left(H_{0,1}^{\intercal}H_{0,2}+H_{0,2}^{\intercal}H_{0,1}\right)\|{\boldsymbol{x}}\|.$
@@
Then, letting $N\coloneqq(k^{\prime}-\deg{F})/2\in\mathbb{N}$, $S_{0}\coloneqq H_{0,1}^{\intercal}H_{0,1}+H_{0,2}^{\intercal}H_{0,2}\|{\boldsymbol{x}}\|^{2}$ and $S\coloneqq H_{1}^{\intercal}H_{1}+H_{2}^{\intercal}H_{2}\|{\boldsymbol{x}}\|^{2}$, we obtain the desired conclusion. ∎
@@
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in $\boldsymbol{x}$} ($\deg S_{0}=N+\deg F$ and $\deg_{{\boldsymbol{x}}}S=N+\deg F-\deg_{{\boldsymbol{x}}}G$), and satisfy that $S_{0}({\boldsymbol{z}}\circ{\boldsymbol{z}})$ (resp., $S({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})$) are SOS matrices in $\boldsymbol{z}$} (resp., $\boldsymbol{z}$} and $\boldsymbol{y}$}), such that
@@
Consider the polynomial matrices $F({\boldsymbol{z}}\circ{\boldsymbol{z}})\in{\mathbb{S}}[{\boldsymbol{z}}]^{p}$ and $G({\boldsymbol{z}}\circ{\boldsymbol{z}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{q}$. Then, by assumption and Theorem 4.1, there exists $N$ and polynomial matrices $H_{0}\in{\mathbb{R}}[{\boldsymbol{z}}]^{\ell_{0}\times p}$, $H\in{\mathbb{R}}[{\boldsymbol{z}},{\boldsymbol{y}}]^{\ell\times pq}$ for some $\ell_{0},\ell\in\mathbb{N}$ that are homogeneous in $\boldsymbol{z}$}, such that
@@
Moreover, by Theorem 4.1, $\deg H_{0}=N+\deg F$, and $\deg_{{\boldsymbol{x}}}H=N+\deg F-\deg_{{\boldsymbol{x}}}G$. Note that there are sets of polynomial matrices
$\{H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}}]^{\ell_{0}\times p}\quad\text{and}\quad\{H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}):{\boldsymbol{\alpha}}\in\{0,1\}^{n}\}\subset{\mathbb{R}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{\ell\times p}$
@@
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})$
holds for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}$. As ${\mathbb{R}}^{n}_{+}$ has interior points, the polynomial matrices on the left and right side of the above equality are identical. Letting
$S_{0}({\boldsymbol{x}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{0,{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}})H_{0,{\boldsymbol{\alpha}}}({\boldsymbol{x}})\text{ and }S({\boldsymbol{x}},{\boldsymbol{y}})\coloneqq\sum_{{\boldsymbol{\alpha}}\in\{0,1\}^{n}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}H_{{\boldsymbol{\alpha}}}^{\intercal}({\boldsymbol{x}},{\boldsymbol{y}})H_{{\boldsymbol{\alpha}}}({\boldsymbol{x}},{\boldsymbol{y}}),$
the conclusion follows. ∎
@@
[42, Theorem 3] Suppose that the polynomial matrix $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is homogeneous and $H({\boldsymbol{x}})\succeq\lambda I_{p}$ for some $\lambda>0$ on $\{{\boldsymbol{x}}\in{\mathbb{R}}_{+}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}$. Write $H({\boldsymbol{x}})=\sum_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}H_{{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ with each $H_{{\boldsymbol{\alpha}}}\in{\mathbb{S}}^{p}$ and let
$L(H)\coloneqq\max_{{\boldsymbol{\alpha}}\in\mathrm{supp}(H)}\frac{{\boldsymbol{\alpha}}!}{\deg H!}\|H_{{\boldsymbol{\alpha}}}\|,$
where $\|\cdot\|$ denotes the spectral norm. Then, for all
$N\geq\frac{\deg H(\deg(H)-1)L(H)}{2\lambda}-\deg H,$
the coefficient of ${\boldsymbol{x}}^{{\boldsymbol{\alpha}}}$ in $({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}H({\boldsymbol{x}})$ for each ${\boldsymbol{\alpha}}\in\mathbb{N}^{n}_{N+\deg H}$ is PD.
We will use this theorem as another intermediary result. Let $D\coloneqq\max\left\{\deg F,\deg_{{\boldsymbol{x}}}G\right\}$.
@@
$F_{\varepsilon}({\boldsymbol{x}})\coloneqq({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg F}F({\boldsymbol{x}})-\varepsilon\left(({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D}I_{p}+\int_{\mathcal{Y}}\left\langle({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{D-\deg_{{\boldsymbol{x}}}G}I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$
@@
By homogeneity, we only need to prove $F_{\varepsilon}\succ 0$ on $\overline{\mathcal{X}}\coloneqq{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\cap\{{\boldsymbol{x}}\in{\mathbb{R}}^{n}\mid{\boldsymbol{e}}^{\intercal}{\boldsymbol{x}}=1\}.$ That is, there exists $\varepsilon>0$ such that $F({\boldsymbol{x}})-\varepsilon\left(I_{p}+\int_{\mathcal{Y}}\left\langle I_{pq},G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}})\right)\succ 0$ for all ${\boldsymbol{x}}\in\overline{\mathcal{X}}$. Since $\overline{\mathcal{X}}$ is compact and $F({\boldsymbol{x}})\succ 0$ on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, letting
@@
we obtain the desired conclusion. ∎
We now prove Theorem 4.2.